OpenStudy (anonymous):

A bus travels 280 km south along a straight path with an average velocity of 88 km/h to the south. The bus stops for 24 min, then it travels 210 km south with an average velocity of 75 km/h to the south. a) How long does the total trip last? b) What is the average velocity for the total trip?

OpenStudy (anonymous):

I got 23011.6s for part A

OpenStudy (anonymous):

And I got 22.6 m/s for part B

OpenStudy (anonymous):

Alright. First off, to calculate the actual time it took for complete the trip: 3.18 h. (190 minutes) to drive 280 km. 24 minute stop time 2.8 h. (168 minutes) to drive 210 km. Total Time: 382 minutes Part B: To calculate the average velocity for the entire trip: v=d/t Simplistic enough. Just convert quickly... 382*60/1=22920 s 490*1000/1=490,000 m. 490,000/22920=21.37 m/s

OpenStudy (anonymous):

to *